Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

A game has 40 chips, each labeled with a number from 1 through 40. Ben randomly chooses one chip. What is the probability that he chooses 10 or a multiple of 8? A.1/8 B.3/20 C.1/6 D.1/320

OpenStudy (anonymous):

B? @peachpi

OpenStudy (anonymous):

@peachpi @LynFran

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5/40 + 1/40 = 6/40 = 3/20

OpenStudy (anonymous):

yeahh. :) thanks @peachpi

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!